OpenStudy (anonymous):

find the indefinite integral. ∫1/3x+2 dx

OpenStudy (kropot72):

Is this it? \[\int\limits_{}^{}(\frac{1}{3}x+2)\ dx\]

OpenStudy (anonymous):

not 1/3

OpenStudy (anonymous):

1 is the only umber in the numerator

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

So it is \[\int\limits_{}^{}(\frac{1}{3x}+2)\ dx\] Correct?

OpenStudy (anonymous):

put the +2 next to the 3x and yu got it lol

OpenStudy (kropot72):

Got it! \[\int\limits_{}^{}\frac{1}{3x+2}\ dx\]

OpenStudy (anonymous):

lol yes! (:

OpenStudy (anonymous):

I think your helper left. This is a simple integral. Most people would use a u-substitution, so I will do so as well. Let u=3x +2, then du =3dx. Solving for dx, we get that \[\frac{ du }{ 3 }=dx\] Now, substituting all of this into your integral, you get: \[\frac{ 1 }{ 3 } \int\limits{\frac{ du }{ u}}=\frac{ 1 }{ 3}\ln \left| u \right| +c\] Now, go back and replace u. We get \[\frac{ 1 }{ 3 }\ln \left| 3x+2 \right|+c\]